# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v2

@@ II.4 Relationship to singular value decomposition
When the effective rank (defined below) is small, the random-mode model “looks like” the SVD at large $N$ since the effects of the $\mathcal{O}(1/\sqrt{N})$ deviations from orthonormality in the sampled modes are negligible. Indeed, a well-known result in high-dimensional geometry is that finite numbers of random vectors with i.i.d. components approximate orthonormal bases at large $N$ [38]. When the effective rank is not small, overlaps between different modes produce a discrepancy between the component strengths $D_{a}$ and singular values $S_{a}$. We now characterize this discrepancy analytically.
We denote the $n$-th moments and participation ratio of the component-strength distribution by
@@
The effective rank is related to the participation ratio of the squared singular value spectrum  [^3] This follows from using Wick’s theorem to evaluate the expectations of $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ and $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ for the numerator and denominator, respectively, of $\text{PR}^{S}$. through
@@
Expanding in small $\alpha\text{PR}^{D}$, $\text{PR}^{S}=\alpha\text{PR}^{D}+\mathcal{O}((\alpha\text{PR}^{D})^{2})$, demonstrating that the component strengths are closely related to the singular values in the low-dimensional regime. While we express the analytic results in this paper in terms of $\alpha\text{PR}^{D}$, the above equation can always be used to translate between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$.
We further quantify this discrepancy in the case where $D_{a}=1$ for all $a$. Using methods from free probability theory [40, 41] (Appendix C), the $M$ singular values are distributed over a range with boundaries
@@
For small $\alpha$, $S_{\pm}=1\pm\sqrt{2\alpha}+\mathcal{O}(\alpha)$. While the random-mode model spreads the nonzero singular values over a range, this spread becomes negligible for small $\alpha$. More generally, the distributions of $D_{a}$ and $S_{a}$ coincide for small $\alpha\text{PR}^{D}$.
@@ III.1 Recurrent neural network model
We study a recurrent neural network of $N$ neurons. Each neuron $i\in\{1,\ldots,N\}$ is characterized by its pre-activation $x_{i}(t)$ and activation $\phi_{i}(t)=\phi(x_{i}(t))$, where $\phi(\cdot)$ is a scalar nonlinearity; we use $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$  [^4] This behaves similarly to the more conventional $\phi(x)=\tanh(x)$ but allows for analytical evaluation of Gaussian integrals.. The network dynamics are governed by
$T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t),\>\>T[x](t)=(1+\partial_{t})x(t),$ (15)
where $J_{ij}$ denotes the synaptic coupling from neuron $j$ to neuron $i$, and $T[\cdot]$ is a causal functional that specifies the single-neuron dynamics. We expect our results to be agnostic to the specific choice of $T[\cdot]$. We next introduce correlation functions that capture single-neuron and collective activity properties.
@@ III.2 Two-point functions and duality of neuronal and temporal covariances
Figure 3: Duality of neuron-by-neuron and time-by-time covariances and its relation to the dimension of activity. Both plots are based on the same simulation of a network of $N=2500$ neurons with $g=2.25$ and i.i.d. couplings. The dimension can be computed either by computing the statistics of the off-diagonals of the neuron-by-neuron covariance $C^{\phi}_{ij}(0)$ (left), or by computing the fluctuations away from the translation-invariant mean-field form of the time-by-time covariance $C^{\phi}(t_{1},t_{2})$ (right).
Let $a\in\{x,\phi\}$ denote either pre-activations or activations. We define two complementary covariance functions, averaging over either time or neurons:
@@
These functions exhibit a form of duality: for any lag $\tau$, the matrix $\bm{C}^{a}(\tau)$ with elements $C^{a}_{ij}(\tau)$ indexed by $(i,j)$ and the matrix with elements $C^{a}(t_{1},t_{2}+\tau)$ indexed by $(t_{1},t_{2})$ have the same eigenvalue spectrum up to scaling (to see this, note that both matrices can be computed from a time-by-neuron activity matrix $\bm{A}$ as $\bm{A}^{T}\bm{A}$ and $\bm{A}\bm{A}^{T}$, respectively, which have identical spectra)  [^5] This duality is reminiscent of ergodicity theorems, which relate temporal and spatial averages, but differs in a variety of ways, including that it holds even for finite time intervals, whereas ergodicity concerns infinite-time limits..
We also define the neuron-averaged response function, describing propagation of infinitesimal perturbations:
$S^{a}(t_{1},t_{2})=\frac{1}{N}\sum_{i=1}^{N}\left.\frac{\delta a_{i}(t_{1})}{\delta I_{j}(t_{2})}\right|_{I_{j}=0},$ (18)
where $I_{j}(t)$ is a source term added to the right-hand side of the network dynamics: $T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t)+I_{i}(t)$. Finally, we defined “stationarized” counterparts:
@@ III.3 Four-point functions, dimension of activity, and principal-component timescales
In contrast to the above two-point functions, which characterize single-neuron activity, four-point functions describe features of collective activity. The same four-point functions were used in [30]. Here, we motivate these functions in more detail and express them in two ways via the duality between neuronal and temporal covariances.
To motivate their definition, consider the dimension of activity, which can be quantified as the participation ratio of the spectrum of the equal-time covariance matrix $\bm{C}^{a}(0)$ [44, 45, 29, 46, 47, 48, 9, 49, 50]. Given the eigenvectors $\bm{v}^{a}_{k}$ and eigenvalues $\lambda^{a}_{k}$ for $k=1,\ldots,N$ (as one would compute as part of principal components analysis), the participation ratio is
@@
An advantageous property of the participation ratio is that it can be expressed as the ratio of the squared-trace and Frobenius norm of $\bm{C}^{a}(0)$, both of which can be further expressed in terms of matrix elements $C_{ij}^{a}(0)$:
$\text{PR}^{a}$ $=\frac{\left(\text{tr}\bm{C}^{a}(0)\right)^{2}}{\lVert\bm{C}(0)\rVert_{F}^{2}}$
$=\frac{C^{a}(0)^{2}}{\frac{1}{N}\sum_{i=1}^{N}C_{ii}(0)^{2}+\frac{1}{N}\sum_{i\neq j}C_{ij}(0)^{2}},$ (21)
where $C^{a}(0)$ is given by Eq. 19. Assuming that diagonal elements are uniform with negligible fluctuations, i.e., $C^{a}_{ii}(0)=C^{a}(0)$ for all $i$ (as occurs, for example, under i.i.d. connectivity [10]), this simplifies to
@@
where, following [30], we define the four-point function
@@
In this paper, we focus on a related function $\Psi^{a}(\bm{\tau})$ that includes the diagonal terms:
@@
The neuron-by-neuron (Eq. 24) and time-by-time (Eq. 25) definitions of $\Psi^{a}(\bm{\tau})$ form the basis of the cavity and path-integral calculations of this function, respectively. Finally, the dimension of activity is given by
@@
which, unlike Eq. 22, holds even when single-neuron variances $C^{a}_{ii}(0)$ are nonuniform across neurons.
To study timescales of collective activity, we consider the principal components of activity:
@@
These principal components are the basis of much of modern analysis of high-dimensional neural data. They all have unit variance due to the $1/\sqrt{\lambda^{a}_{k}}$ normalization, but potentially very different characteristic timescales (Fig. 9a). To extract the timescales of just the leading components, we weight them by the squares of their corresponding eigenvalues. This gives:
@@
showing that $\Psi^{a}(\bm{\tau})$ captures the temporal structure of the leading principal components of activity (Fig. 9c,d). Note that weighting by the eigenvalues themselves, rather than their squares, gives $N^{-1}\sum_{k}\lambda^{a}_{k}\left\langle p^{a}_{k}(t)p^{a}_{k}(t+\tau)\right\rangle_{t}=C^{a}(\tau)$, recovering single-neuron information (Fig. 9b,d).
@@ IV Path-integral analysis of the random-mode model
These values also correspond to the saddle point of a path integral. The summary statistics we have defined are self-averaging, meaning that the same values are obtained, up to $\mathcal{O}(1/\sqrt{N})$ fluctuations, when they are computed based on a single realization of a large network. In this paper, we are interested in statistically stationary states such that, for $N\to\infty$, the two-point functions depend on only time differences and thus can be replaced by their stationarized counterparts (Eq. 19), i.e., for all $t$,
@@
Calculating two-point functions is well established “standard DMFT.” As we review in Appendix E, four-point functions were previously calculated under i.i.d. connectivity using a two-site version of the cavity method [30]. In this section, we introduce a new path-integral approach to calculating four-point functions that allows for rapid solution of the i.i.d. model and, importantly, more complicated models including the random-mode model.
@@ IV.1 Relating $\Psi^{a}$ to temporal covariance fluctuations
To cast the problem in the language of path integrals, we turn to the time-by-time definition of $\Psi^{a}(\bm{\tau})$ (Eq. 25). If the time-by-time covariance $C^{a}(t_{1},t_{2})$ were equal to its limiting value, $C^{a}_{\star}(t_{2}-t_{1})$, we would have $\Psi_{\star}^{a}(\bm{\tau})\overset{?}{=}N\left\langle C^{a}_{\star}(t_{2}-t_{1})C^{a}_{\star}(t_{2}-t_{1}+\tau_{2}-\tau_{1})\right\rangle_{t_{1},t_{2}}=0$} since, for large time differences, $C^{a}_{\star}(t_{2}-t_{1})$, decays exponentially in $|t_{2}-t_{1}|$ [10]. However, based on the neuron-by-neuron definition (Eq. 24), $\Psi_{\star}^{a}(\bm{\tau})$ is clearly a nonzero, order-one quantity. This apparent contradiction is resolved by considering the $\mathcal{O}(1/\sqrt{N})$ fluctuations around the limiting value. The necessity of these fluctuations becomes clear when considering the duality between neuronal and temporal covariances: the presence of nonzero cross-covariances $C^{a}_{ij}(0)$ indicates low-dimensional structure in the system; equivalently, this structure manifests as the system revisiting similar states over time, resulting in nonzero temporal covariance $C^{a}(t_{1},t_{2})$ even for large time differences $|t_{2}-t_{1}|\gg 1$ (Fig. 3). To capture this behavior, we define fluctuations around the saddle point as
@@
noting that terms involving $C^{a}_{\star}(t_{2}-t_{1})$ vanish at large time differences. Our task, therefore, is to compute the covariance of the fluctuations, $\delta C^{a}(t_{1},t_{2})$, captured asymtotically by
@@
Then, to compute $\Psi_{\star}^{a}(\bm{\tau})$, we take in Eq. 33 the limits
@@ IV.2 Calculation of two- and four-point functions
Details of our calculation are given in Appendix F. Briefly, we compute the covariance of fluctuations (Eq. 33) using the Martin-Siggia-Rose-De Dominicis-Janssen (MSRDJ) path-integral formalism [51, 52, 53, 54, 55, 56]. For any fixed coupling matrix $\bm{J}$, the path integral is
@@
The auxiliary field $\hat{x}_{i}(t)$ is conjugate to $x_{i}(t)$ and enforces the network’s equations of motion (Eq. 15) through the delta-function representation $(2\pi)^{-1}\int d\hat{x}\,e^{i\hat{x}x}=\delta(x)$. Upon introduction of source terms that are not needed for our purposes, this path integral serves as a generating functional for correlation and response functions.
In principle, to compute $\Psi^{a}_{\star}(\bm{\tau})$ in this framework, after integrating out $\bm{J}$ from $Z[\bm{J}]$ to obtain a statistical field theory governing two-point functions, we (1) find the saddle-point solution for two-point functions by extremizing the action (i.e., standard DMFT); (2) compute the time2-by-time2 Hessian matrix of the action describing fluctuations around this saddle point; (3) invert this Hessian and extract the sub-block corresponding to the covariance of $\delta C^{\phi}(t_{1},t_{2})$ (Eq. 33); and (4) apply the temporal limits (Eq. 34) to this sub-block to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$.
Direct inversion of the time2-by-time2 Hessian is analytically intractable. However, we show that the temporal limits commute with Hessian inversion, allowing us to invert a low-dimensional frequency-dependent Hessian instead (Appendix G). This readily yields the Fourier-space function $\Psi^{\phi}_{\star}(\bm{\omega})$, where $\bm{\omega}=(\omega_{1},\omega_{2})$. A few additional steps yield $\Psi^{x}_{\star}(\bm{\omega})$. This approach allows for the efficient calculation of two and four point functions for a given model, requiring only a few lines of computation directly from the action of the theory.
As a demonstration of this approach, we apply it to a generalization of the i.i.d. model with correlated reciprocal couplings, recovering results of [30] (Appendix H).

### v3

@@ II.4 Relationship to singular-value decomposition
When the effective rank (defined below) is small and $N$ is large, the component strengths and modes approximate singular values and vectors, respectively, since the effects of the $\mathcal{O}(1/\sqrt{N})$ deviations from orthonormality in the sampled modes are negligible. Indeed, a well-known result in high-dimensional geometry is that finite numbers of random vectors with i.i.d. components approximate orthonormal bases at large $N$ [47]. When the effective rank is not small, overlaps between different modes produce a discrepancy between the component strengths $D_{a}$ and singular values $S_{a}$. We now characterize this discrepancy analytically.
We denote the $n$th moments and participation ratio of the component-strength distribution by
@@
The effective rank is related to the participation ratio of the squared singular-value spectrum at large $N$  [^3] This follows from using Wick’s theorem to evaluate the expectations of $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ and $\text{tr}\left(\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\bm{L}\bm{D}\bm{R}^{T}\bm{R}\bm{D}\bm{L}^{T}\right)$ for the numerator and denominator, respectively, of $\text{PR}^{S}$. through
@@
Expanding in small $\alpha\text{PR}^{D}$, $\text{PR}^{S}=\alpha\text{PR}^{D}+\mathcal{O}\bm{(}(\alpha\text{PR}^{D})^{2}\bm{)}$, demonstrating that the component strengths are closely related to the singular values in the low dimensional regime. While we express the analytic results in this paper in terms of $\alpha\text{PR}^{D}$, the above equation can always be used to translate between $\alpha\text{PR}^{D}$ and $\text{PR}^{S}$.
We further quantify this discrepancy in the case where $D_{a}=1$ for all $a$. Using methods from free probability theory [49, 50] (Appendix C; Fig. 8), the $M$ singular values are distributed over a range with boundaries
@@
For small $\alpha$, $S_{\pm}=1\pm\sqrt{2\alpha}+\mathcal{O}(\alpha)$. Thus, while the random-mode model spreads the nonzero singular values over a range, this spread becomes negligible for small $\alpha$. More generally, the distributions of $D_{a}$ and $S_{a}$ coincide for small $\alpha\text{PR}^{D}$.
@@ III.1 Recurrent neural network model
We study a recurrent neural network of $N$ neurons. Each neuron $i\in\{1,\ldots,N\}$ is characterized by its preactivation $x_{i}(t)$ and activation $\phi_{i}(t)=\phi\bm{(}x_{i}(t)\bm{)}$, where $\phi(\cdot)$ is a scalar nonlinearity; we use $\phi(x)=\text{erf}\left({\sqrt{\pi}x}/{2}\right)$, which is sigmoid-shaped and has $\phi^{\prime}(0)=1$  [^4] This behaves similarly to the more conventional $\phi(x)=\tanh(x)$ but allows for analytical evaluation of Gaussian integrals.. The network dynamics are governed by
$T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t),$ (15a)
$\text{where}\quad T[x](t)=(1+\partial_{t})x(t).$ (15b)
Here, $J_{ij}$ denotes the synaptic coupling from neuron $j$ to neuron $i$, and $T[\cdot]$ is a causal functional that specifies the single-neuron dynamics for which Eq. (15b) is a canonical choice. We expect our results to be agnostic to the specific choice of $T[\cdot]$. We next introduce correlation functions that capture single-neuron and collective activity properties.
@@ III.2 Two-point functions and duality of neuronal and temporal covariances
Figure 3: Duality of neuron-by-neuron and time-by-time covariances and its relation to the dimension of activity. Both plots are based on the same simulation of a network of $N=2500$ neurons with $g=2.25$ and i.i.d. couplings. The dimension can be computed either by computing the statistics of the off diagonals of the neuron-by-neuron covariance $C^{\phi}_{ij}(0)$ (left), or by computing the fluctuations away from the translation-invariant mean-field form of the time-by-time covariance $C^{\phi}(t_{1},t_{2})$ (right).
Let $a\in\{x,\phi\}$ denote either preactivations or activations. We define two complementary covariance functions, averaging over either time or neurons,
@@
These functions exhibit a form of duality: for any lag $\tau$, the matrix $\bm{C}^{a}(\tau)$ with elements $C^{a}_{ij}(\tau)$ indexed by $(i,j)$ and the matrix with elements $C^{a}(t_{1},t_{2}+\tau)$ indexed by $(t_{1},t_{2})$ have the same eigenvalue spectrum, up to scaling (to see this, note that both matrices can be computed from a time-by-neuron activity matrix $\bm{A}$ as $\bm{A}^{T}\bm{A}$ and $\bm{A}\bm{A}^{T}$, respectively, which have identical spectra)  [^5] This duality is reminiscent of ergodicity theorems, which relate temporal and spatial averages, but differs in a variety of ways, including that it holds even for finite-time intervals, whereas ergodicity concerns infinite-time limits..
We also define the neuron-averaged response function, describing the propagation of infinitesimal perturbations:
$S^{a}(t_{1},t_{2})=\frac{1}{N}\sum_{i=1}^{N}\left.\frac{\delta a_{i}(t_{1})}{\delta I_{i}(t_{2})}\right|_{\bm{I}=0},$ (18)
where $I_{i}(t)$ is a source term added to the right-hand side of the network dynamics: $T[x_{i}](t)=\sum_{j=1}^{N}J_{ij}\phi_{j}(t)+I_{i}(t)$. Finally, we define “stationarized” counterparts,
@@ III.3 Four-point functions, dimension of activity, and principal-component timescales
In contrast to the above two-point functions, which characterize single-neuron activity, four-point functions describe features of collective activity. The same four-point functions were used in Ref. [39]. Here, we motivate these functions in more detail and express them in two ways via the duality between neuronal and temporal covariances.
To motivate their definition, consider the dimension of activity, which can be quantified as the participation ratio of the spectrum of the equal-time covariance matrix $\bm{C}^{a}(0)$ [53, 54, 36, 55, 56, 57, 9, 58, 59]. Given the eigenvectors $\bm{v}^{a}_{k}$ and eigenvalues $\lambda^{a}_{k}$ for $k=1,\ldots,N$, as one would compute as part of principal components analysis, the participation ratio is
@@
An advantageous property of the participation ratio is that it can be expressed as the ratio of the squared trace and Frobenius norm of $\bm{C}^{a}(0)$, both of which can be further expressed in terms of matrix elements $C_{ij}^{a}(0)$,
$\text{PR}^{a}$ $=\frac{\left(\text{tr}\bm{C}^{a}(0)\right)^{2}}{\lVert\bm{C}^{a}(0)\rVert_{F}^{2}}$
$=\frac{C^{a}(0)^{2}}{\frac{1}{N}\sum_{i=1}^{N}C^{a}_{ii}(0)^{2}+\frac{1}{N}\sum_{i\neq j}C^{a}_{ij}(0)^{2}},$ (21)
where $C^{a}(0)$ is given by Eq. (19). Assuming that diagonal elements are uniform with negligible fluctuations, i.e., $C^{a}_{ii}(0)=C^{a}(0)$ for all $i$ (as occurs, for example, under i.i.d. connectivity [10]), this simplifies to
@@
where, following Ref. [39], we define the four-point function
@@
In this paper, we focus on a related function $\Psi^{a}(\bm{\tau})$ that includes the diagonal terms,
@@
The neuron-by-neuron [Eq. (24)] and time-by-time [Eq. (25)] definitions of $\Psi^{a}(\bm{\tau})$ form the basis of the cavity and path-integral calculations of this function, respectively. Finally, the dimension of activity is given by
@@
which, unlike Eq. (22), holds even when single-neuron variances $C^{a}_{ii}(0)$ are nonuniform across neurons.
To study timescales of collective activity, we consider the principal components of activity,
@@
These principal components are the basis of much of modern analysis of high-dimensional neural data [35, 36, 37, 38]. They all have unit variance due to the $1/\sqrt{\lambda^{a}_{k}}$ normalization, but potentially very different characteristic timescales [Fig. 9(a)]. To extract the timescales of just the leading components, we weight them by the squares of their corresponding eigenvalues. This gives
@@
showing that $\Psi^{a}(\bm{\tau})$ captures the temporal structure of the leading principal components of activity [Figs. 9(c) and 9(d)]. Note that weighting by the eigenvalues themselves, rather than their squares, gives $N^{-1}\sum_{k}\lambda^{a}_{k}\left\langle p^{a}_{k}(t)p^{a}_{k}(t+\tau)\right\rangle_{t}=C^{a}(\tau)$, recovering single-neuron information [Figs. 9(b) and 9(d)].
@@ IV Path-integral analysis of the random-mode model
These values also correspond to the saddle point of a path integral. The summary statistics we have defined are self-averaging, meaning that the same values are obtained, up to $\mathcal{O}(1/\sqrt{N})$ fluctuations, when they are computed based on a single realization of a large network. In this paper, we are interested in statistically stationary states such that, for $N\to\infty$, the two-point functions depend on time differences only and thus can be replaced by their stationarized counterparts [Eq. (19)], i.e., for all $t$,
@@
Calculating two-point functions is well-established “standard DMFT” [60, 61, 62]. As we review in Appendix E, four-point functions were previously calculated for networks with i.i.d. couplings using a two-site version of the cavity method [39]. In this section, we introduce a new path-integral approach to calculating four-point functions that allows for the rapid solution of the i.i.d. model and generalizes to more complicated models, including the random-mode model.
@@ IV.1 Relating $\Psi^{a}$ to temporal covariance fluctuations
To cast the problem in the language of path integrals, we turn to the time-by-time definition of $\Psi^{a}(\bm{\tau})$ [Eq. (25)]. If the time-by-time covariance $C^{a}(t_{1},t_{2})$ were equal to its limiting value, $C^{a}_{\star}(t_{2}-t_{1})$, we would have
$\Psi_{\star}^{a}(\bm{\tau})\overset{?}{=}N\left\langle C^{a}_{\star}(t_{2}-t_{1})C^{a}_{\star}(t_{2}-t_{1}+\tau_{2}-\tau_{1})\right\rangle_{t_{1},t_{2}}=0$} (31)
since, for large time differences, $C^{a}_{\star}(t_{2}-t_{1})$ decays exponentially in $|t_{2}-t_{1}|$ [10]. However, based on the neuron-by-neuron definition [Eq. (24)], $\Psi_{\star}^{a}(\bm{\tau})$ is clearly a nonzero, order-one quantity. This apparent contradiction is resolved by considering the $\mathcal{O}(1/\sqrt{N})$ fluctuations around the limiting value. The necessity of these fluctuations becomes clear when considering the duality between neuronal and temporal covariances: the presence of nonzero cross-covariances $C^{a}_{ij}(0)$ indicates low dimensional structure in the system; equivalently, this structure manifests as the system revisiting similar states over time, resulting in nonzero temporal covariance $C^{a}(t_{1},t_{2})$ even for large time differences $|t_{2}-t_{1}|\gg 1$ (Fig. 3). To capture this behavior, we define fluctuations around the saddle point as
@@
noting that terms involving $C^{a}_{\star}(t_{2}-t_{1})$ vanish at large time differences. Our task, therefore, is to compute the covariance of the fluctuations, $\delta C^{a}(t_{1},t_{2})$, captured asymptotically by
@@
Then, to compute $\Psi_{\star}^{a}(\bm{\tau})$, we take in Eq. (34) the limits
@@ IV.2 Calculation of two- and four-point functions
Details of the calculation are given in Appendix F. Briefly, we compute the covariance of fluctuations [Eq. (34)] using the Martin-Siggia-Rose-Janssen–de Dominicis (MSRJD) path-integral formalism [63, 64, 65, 60, 61, 62]. For any fixed coupling matrix $\bm{J}$, the path integral is
@@
The auxiliary field $\hat{x}_{i}(t)$ is conjugate to $x_{i}(t)$ and enforces the network’s equations of motion [Eq. (15a)] through the $\delta$-function representation $(2\pi)^{-1}\int d\hat{x}\,e^{i\hat{x}x}=\delta(x)$. Upon introduction of source terms (that are not needed for our purposes), this path integral serves as a generating functional for correlation and response functions.
In principle, to compute $\Psi^{a}_{\star}(\bm{\tau})$ in this framework, after integrating out $\bm{J}$ from $Z[\bm{J}]$ to obtain a statistical field theory governing two-point functions, we (1) find the saddle-point solution for two-point functions by extremizing the action (i.e., standard DMFT); (2) compute the time2-by-time2 Hessian matrix of the action describing fluctuations around this saddle point; (3) invert this Hessian and extract the sub-block corresponding to the covariance of $\delta C^{\phi}(t_{1},t_{2})$ [Eq. (34)]; and (4) apply the temporal limits [Eq. (35)] to this sub-block to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$.
Direct inversion of the time2-by-time2 Hessian is analytically intractable. However, we show that the temporal limits commute with Hessian inversion, allowing us to invert a low dimensional frequency-dependent Hessian instead (Appendix G). This readily yields the Fourier-space function $\Psi^{\phi}_{\star}(\bm{\omega})$, where $\bm{\omega}=(\omega_{1},\omega_{2})$. A few additional steps yield $\Psi^{x}_{\star}(\bm{\omega})$. This approach allows for the efficient calculation of two- and four-point functions for a given model, requiring only a few lines of computation directly from the action of the theory.
As a demonstration of this approach, we apply it to a generalization of the i.i.d. model with correlated reciprocal couplings, recovering results of Ref. [39] (Appendix H).